OpenStudy (anonymous):

-3 (6x - 4) + 11x = -9

OpenStudy (anonymous):

-18x+12+11x=-9 -7x+12=-9 -7x=-21 x=3

sam (.sam.):

Multiply -3 by each term inside the parentheses -18x+12+11x=-9 Since -18x and 11x are like terms, subtract 11x from -18x to get -7x. -7x+12=-9 Move all terms not containing x to the right-hand side of the equation. -7x=-21 Try to do from here?

OpenStudy (anonymous):

thx both but idk who to give the medal

sam (.sam.):

@Outkast3r09 please don't give full answers without explaining

OpenStudy (anonymous):

Give it to him I don't care about the medals . I simply look for if one has been given out so I know the question has been answered..

OpenStudy (anonymous):

ow

sam (.sam.):

Even though you don't care about medals its against the rules

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

What are you even going on about? Listen, I'm not hear to pick fights with other people so I'm just going to end this. The only thing I said was that I don't care if I get the medal because I don't believe that "medals" are worth more than helping people I gave just as good as any Solution manual would give.

OpenStudy (anonymous):

ya i agre

OpenStudy (farmdawgnation):

Hey @Outkast3r09 - Sam is referring to our code of conduct which can be found here: http://openstudy.com/code-of-conduct As we state in the CoC, we as that our users "encourage and guide those needing help, and not just give them an answer". While your answer falls in a grey zone of sorts because you are providing the supporting work behind your answer, I'd encourage you to try and engage the asker to work at least part of the problem on their own. Thanks for your contributions to OpenStudy.